# Adaptive Testing for LLM Evaluation: A Psychometric Alternative to Static Benchmarks

### v1

@@ Appendix G Additional Experimental Results
Figure 5: Comparison of raw average scores and whole-bank ability estimates on WinoGrande. (Left) Whole-bank estimates show a non-linear relationship with average score and reveal clearer separation on high-performing models, highlighting that ability captures relative item difficulty and provides finer differentiation beyond raw accuracy. (Right) Rank comparison indicates strong but imperfect alignment (Spearman $\rho=0.86$, Kendall $\tau=0.70$), with deviations from the diagonal reflecting cases where ability-based ranking distinguishes models more effectively than accuracy alone.
Figure 6: Comparison of raw average scores and whole-bank ability estimates on ARC. (Left) Whole-bank estimates exhibit a non-linear relationship with average scores, providing clearer separation on high-performing models by accounting for item difficulty and discrimination. (Right) Rank comparison shows strong but not perfect alignment between the two metrics (Spearman $\rho=0.91$, Kendall $\tau=0.77$), with deviations from the diagonal highlighting cases where ability-based ranking offers more informative distinctions than raw accuracy alone.
Figure 7: Two models with the similar average accuracy (0.713) and (0.714) on ARC nevertheless receive very different whole-bank ability estimates. Model A (mera-mix-4x7B) attains a whole-bank ability rank of 270 because its correct responses are concentrated on more difficult items. In contrast, Model B (LLaMAAntino-3-ANITA-8B-Inst-DPO-ITA) is assigned a much lower whole-bank ability rank of 2612, as its successes occur primarily on easier items. This divergence shows how IRT-based ability estimation can distinguish models that appear identical under raw accuracy by accounting for item difficulty.
Figure 8: Two models with the same average accuracy (0.853) on HellaSwag nevertheless receive very different whole-bank ability estimates. Model A (supermario_v1) attains a whole-bank ability rank of 347 because its correct responses are concentrated on more difficult items. In contrast, Model B (contaminated_proof_7b_v1.0_safetensor) is assigned a much lower whole-bank ability rank of 3074, as its successes occur primarily on easier items. This divergence shows how IRT-based ability estimation can distinguish models that appear identical under raw accuracy by accounting for item difficulty.

### v2

@@ G.3 Experiment on MMLU
MMLU-High School Physics (119 items) ATLASSE≤0.1 0.026 63.30 0.5825 0.5317 3.54
ATLASSE≤0.2 0.055 51.75 0.4900 0.4356 3.08
ATLASSE≤0.3 0.106 20.30 0.3393 0.1932 1.37
MMLU-High School Psychology (285 items) ATLASSE≤0.1 0.035 100.00 0.6086 0.3509 6.19
ATLASSE≤0.2 0.147 26.80 0.2433 0.1284 1.97
ATLASSE≤0.3 0.196 11.48 0.2101 0.0881 0.92
MMLU-High School Statistics (185 items) ATLASSE≤0.1 0.038 69.47 0.5587 0.3757 4.07
ATLASSE≤0.2 0.154 21.48 0.2987 0.1219 1.67
ATLASSE≤0.3 0.205 11.81 0.2483 0.0700 1.03
MMLU-High School US History (183 items) ATLASSE≤0.1 0.030 96.68 0.7761 0.5283 5.77
ATLASSE≤0.2 0.069 73.31 0.6137 0.4027 4.80
ATLASSE≤0.3 0.151 13.62 0.3348 0.1319 1.05
MMLU-High School World History (206 items) ATLASSE≤0.1 0.031 96.43 0.7407 0.4680 5.88
ATLASSE≤0.2 0.108 40.83 0.3806 0.2580 2.85
ATLASSE≤0.3 0.176 17.10 0.2629 0.1230 1.28
MMLU-Human Aging (188 items) ATLASSE≤0.1 0.034 100.00 0.7341 0.5319 5.86
ATLASSE≤0.2 0.108 40.24 0.3546 0.2515 2.69
ATLASSE≤0.3 0.154 27.38 0.2445 0.1802 2.00
MMLU-Human Sexuality (116 items) ATLASSE≤0.1 0.032 100.00 0.9137 0.8621 6.07
ATLASSE≤0.2 0.062 83.66 0.7655 0.7212 5.25
ATLASSE≤0.3 0.154 20.43 0.2825 0.1759 1.51
MMLU-International Law (103 items) ATLASSE≤0.1 0.028 100.00 0.9780 0.9709 5.88
ATLASSE≤0.2 0.061 76.17 0.7582 0.7400 4.72
ATLASSE≤0.3 0.105 34.28 0.4298 0.3333 2.57
MMLU-Jurisprudence (99 items) ATLASSE≤0.1 0.021 99.00 1.0000 1.0000 0.79
ATLASSE≤0.2 0.047 71.20 0.7262 0.7191 0.60
ATLASSE≤0.3 0.095 33.57 0.3986 0.3399 0.38
MMLU-Logical Fallacies (147 items) ATLASSE≤0.1 0.030 100.00 0.8562 0.6803 5.92
ATLASSE≤0.2 0.064 70.48 0.6483 0.4928 4.73
ATLASSE≤0.3 0.117 22.21 0.2938 0.1632 1.69
MMLU-Machine Learning (98 items) ATLASSE≤0.1 0.027 74.29 0.7619 0.7586 3.34
ATLASSE≤0.2 0.069 39.44 0.4402 0.4018 1.89
ATLASSE≤0.3 0.143 13.81 0.3032 0.1410 0.71
MMLU-Management (92 items) ATLASSE≤0.1 0.022 92.00 1.0000 1.0000 0.74
ATLASSE≤0.2 0.024 90.39 0.9831 0.9831 0.72
ATLASSE≤0.3 0.059 44.91 0.5238 0.4881 0.47
MMLU-Marketing (178 items) ATLASSE≤0.1 0.033 100.00 0.7974 0.5618 5.78
ATLASSE≤0.2 0.081 61.18 0.5305 0.3645 4.06
ATLASSE≤0.3 0.134 25.09 0.3035 0.1660 1.84
MMLU-Medical Genetics (87 items) ATLASSE≤0.1 0.024 84.38 0.9699 0.9698 0.74
ATLASSE≤0.2 0.045 69.31 0.7995 0.7967 0.62
ATLASSE≤0.3 0.072 55.16 0.6449 0.6349 0.51
MMLU-Miscellaneous (314 items) ATLASSE≤0.1 0.048 65.51 0.3154 0.2085 3.66
ATLASSE≤0.2 0.156 24.00 0.1576 0.0991 1.58
ATLASSE≤0.3 0.203 15.73 0.1295 0.0697 1.10
MMLU-Moral Disputes (229 items) ATLASSE≤0.1 0.041 99.73 0.6574 0.4355 6.03
ATLASSE≤0.2 0.142 27.23 0.2498 0.1526 1.94
ATLASSE≤0.3 0.195 13.45 0.2077 0.1090 0.98
MMLU-Moral Scenarios (486 items) ATLASSE≤0.1 0.059 34.18 0.1610 0.0704 2.24
ATLASSE≤0.2 0.114 19.23 0.1216 0.0436 1.41
ATLASSE≤0.3 0.158 14.14 0.0986 0.0334 1.05
MMLU-Nutrition (244 items) ATLASSE≤0.1 0.039 95.50 0.6325 0.3914 6.00
ATLASSE≤0.2 0.129 38.88 0.3028 0.1953 2.89
ATLASSE≤0.3 0.176 21.25 0.2221 0.1191 1.69
MMLU-Philosophy (214 items) ATLASSE≤0.1 0.041 96.45 0.6360 0.4510 5.90
ATLASSE≤0.2 0.152 23.09 0.1983 0.1258 1.89
ATLASSE≤0.3 0.202 16.56 0.1657 0.0927 1.37
MMLU-Prehistory (265 items) ATLASSE≤0.1 0.039 93.64 0.6004 0.3531 5.72
ATLASSE≤0.2 0.099 44.74 0.4126 0.2078 3.28
ATLASSE≤0.3 0.190 12.29 0.2773 0.0989 0.95
MMLU-Professional Accounting (221 items) ATLASSE≤0.1 0.047 63.18 0.4458 0.2861 3.83
ATLASSE≤0.2 0.170 14.16 0.2016 0.0682 1.03
ATLASSE≤0.3 0.220 10.46 0.1899 0.0647 0.83
MMLU-Professional Law (518 items) ATLASSE≤0.1 0.064 26.94 0.1811 0.0519 1.84
ATLASSE≤0.2 0.151 10.88 0.1380 0.0283 0.98
ATLASSE≤0.3 0.188 10.06 0.1308 0.0271 0.90
MMLU-Professional Medicine (218 items) ATLASSE≤0.1 0.044 95.14 0.6919 0.4363 6.11
ATLASSE≤0.2 0.167 24.91 0.3403 0.1649 2.12
ATLASSE≤0.3 0.220 10.25 0.2794 0.1019 0.95
MMLU-Professional Psychology (344 items) ATLASSE≤0.1 0.051 58.00 0.3224 0.1685 3.86
ATLASSE≤0.2 0.166 11.53 0.1850 0.0670 1.06
ATLASSE≤0.3 0.205 10.08 0.1715 0.0645 0.95
MMLU-Public Relations (93 items) ATLASSE≤0.1 0.027 93.05 1.0000 1.0000 0.74
ATLASSE≤0.2 0.048 72.97 0.7926 0.7854 0.61
ATLASSE≤0.3 0.095 24.50 0.3729 0.2633 0.31
MMLU-Security Studies (201 items) ATLASSE≤0.1 0.038 100.00 0.7247 0.4975 6.11
ATLASSE≤0.2 0.119 34.74 0.3103 0.2038 2.59
ATLASSE≤0.3 0.171 17.12 0.1992 0.1103 1.30
MMLU-Sociology (166 items) ATLASSE≤0.1 0.032 100.00 0.8412 0.6024 6.02
ATLASSE≤0.2 0.070 72.92 0.6391 0.5095 5.00
ATLASSE≤0.3 0.137 37.22 0.3946 0.2712 2.70
MMLU-US Foreign Policy (90 items) ATLASSE≤0.1 0.024 90.00 1.0000 1.0000 0.72
ATLASSE≤0.2 0.026 89.55 0.9950 0.9950 0.71
ATLASSE≤0.3 0.063 30.48 0.4204 0.3389 0.37
MMLU-Virology (140 items) ATLASSE≤0.1 0.032 100.00 0.8644 0.7143 6.01
ATLASSE≤0.2 0.081 63.64 0.5787 0.4711 4.40
ATLASSE≤0.3 0.134 24.11 0.2965 0.1926 1.84
MMLU-World Religions (137 items) ATLASSE≤0.1 0.033 100.00 0.8796 0.7299 6.05
ATLASSE≤0.2 0.068 87.19 0.7734 0.6657 5.23
ATLASSE≤0.3 0.134 37.56 0.4063 0.2984 2.68
TruthfulQA
@@
WinoGrande
Figure 5: Comparison of raw average scores and whole-bank ability estimates on WinoGrande. (Left) Whole-bank ability estimates show a non-linear relationship with average score and reveal clearer separation on high-performing models, highlighting that ability captures relative item difficulty and provides finer differentiation beyond raw accuracy. (Right) Rank comparison indicates strong but imperfect alignment (Spearman $\rho=0.86$, Kendall $\tau=0.70$), with deviations from the diagonal reflecting cases where ability-based ranking distinguishes models more effectively than accuracy alone.
ARC
Figure 6: Comparison of raw average scores and whole-bank ability estimates on ARC. (Left) Whole-bank ability estimates exhibit a non-linear relationship with average scores, providing clearer separation on high-performing models by accounting for item difficulty and discrimination. (Right) Rank comparison shows strong but not perfect alignment between the two metrics (Spearman $\rho=0.91$, Kendall $\tau=0.77$), with deviations from the diagonal highlighting cases where ability-based ranking offers more informative distinctions than raw accuracy alone.
ARC
Figure 7: Two models with the similar average accuracy (0.713) and (0.714) on ARC nevertheless receive very different whole-bank ability estimates. Difficulty levels are ordered from easiest (D1) to hardest (D10). Model A (mera-mix-4x7B) attains a whole-bank ability rank of 270 because its correct responses are concentrated on more difficult items. In contrast, Model B (LLaMAAntino-3-ANITA-8B-Inst-DPO-ITA) is assigned a much lower whole-bank ability rank of 2612, as its successes occur primarily on easier items. This divergence shows how IRT-based ability estimation can distinguish models that appear identical under raw accuracy by accounting for item difficulty.
HellaSwag
Figure 8: Two models with the same average accuracy (0.853) on HellaSwag nevertheless receive very different whole-bank ability estimates. Difficulty levels are ordered from easiest (D1) to hardest (D10). Model A (supermario_v1) attains a whole-bank ability rank of 347 because its correct responses are concentrated on more difficult items. In contrast, Model B (contaminated_proof_7b_v1.0_safetensor) is assigned a much lower whole-bank ability rank of 3074, as its successes occur primarily on easier items. This divergence shows how IRT-based ability estimation can distinguish models that appear identical under raw accuracy by accounting for item difficulty.